OpenStudy (vortish):

x^2 + 10+21=0 solve the following using the quadratic formula needs to be exact using radicals as needed

OpenStudy (wolf1728):

Is that typed correctly? There seems to be no 'x' term in there.

OpenStudy (vortish):

yes it is typed correctly

OpenStudy (wolf1728):

So it is x^2 + 21 = 0 correct?

OpenStudy (vortish):

no x^2 +10 +21=0

OpenStudy (vortish):

x^2 +10x +21=0 my mistake

OpenStudy (wolf1728):

Uh-huh I thought so x^2 + 10x + 21 = 0 a = 1 b=10 c=21 x=-10 +-sqroot(100 - 84) / 2 x = (-10 +sqroot(16))/2 x = -10 + 4 /2 x = -3 x = (-10 -sqroot(16))/2 x = (-10 -4)/2 x = -7

OpenStudy (wolf1728):

and here is the quadratic formula: